OpenStudy (anonymous):

What are the x-intercepts of f(x)= -16x^2+60x+16 ??!

OpenStudy (scoobydo):

X=1

OpenStudy (anonymous):

How'd you get that ?

OpenStudy (anonymous):

@scoobydo

OpenStudy (scoobydo):

without the total u could have X whatever you want (is there choices)

OpenStudy (anonymous):

No. It gives me the expression and says .b" what are the x-intercepts of the graph f(x)? show your work ".

OpenStudy (anonymous):

I'm just trying to learn, and get a good grade.

OpenStudy (scoobydo):

1*1=1 16*1=16 60*1=60 so that means, 1+16+60+16=93

OpenStudy (welshfella):

first simplify it a bit by divding through by -4 = -4(4x^2 - 15x - 16) now we need to solve 4x^2 - 15x - 16 = 0 this will factor

OpenStudy (welshfella):

** sorry that should be 4x^2 - 15x - 4 = 0

OpenStudy (welshfella):

4x^2 - 16x + x - 4 = 0 can you continue from here?

OpenStudy (welshfella):

Oh - I'm talking to myself!!

OpenStudy (scoobydo):

no can u help me please

OpenStudy (johnweldon1993):

Lol but it was so informative for yourself Welsh!

OpenStudy (welshfella):

yea lol

OpenStudy (welshfella):

factor in groups of 2 4x(x - 4) + 1(x - 4) = 0 x - 4 is common so (4x + 1(x - 4) = 0

OpenStudy (welshfella):

now 4x + 1 = 0 or x - 4 = 0 now its easy to find the solution set

OpenStudy (scoobydo):

@welshfella please help me in earth science

OpenStudy (welshfella):

i doubt i can Dont know much about that.